Circular No. 7053

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 W1 (SPAHR) The following CCD astrometric positions have been reported:

1998 UT α_{2000} δ_{2000} m_{1} Observer
Nov. 17.91109 4^{h}32^{m}18.^{s}15 −10^{o}25′02′′.0  Manca
17.96348 4 32 15.62 −10 24 36.0 16  Tichý
17.96637 4 32 15.46 −10 24 34.4  _{”}
17.98176 4 32 14.80 −10 24 26.8  Mikuž
18.04478 4 32 11.80 −10 23 56.6  _{”}
18.17402 4 32 06.56 −10 22 51.6 16.1 Kowalski
18.18895 4 32 05.92 −10 22 44.8 15.9  _{”}
18.29858 4 32 00.83 −10 21 50.5  Spahr
18.47045 4 31 52.60 −10 20 24.8  _{”}
18.47565 4 31 53.30 −10 20 23.4  Kojima
18.49824 4 31 52.19 −10 20 11.8 15.7  _{”}

F. Manca, A. Testa and V. Giuliani (Sormano). 0.5-m reflector. M. Tichý and Z. Moravec (Kleť). 0.57-m f/5.2 reflector. H. Mikuž and S. Matičić (Crni vrh). 0.36-m f/6.7 Schmidt-Cassegrain. R. Kowalski (Zephyrhills). 0.18-m f/9.4 reflector. Coma < 1′. T. B. Spahr (Catalina Sky Survey). 0.41-m f/3 Schmidt. T. Kojima (Chiyoda). 0.25-m f/6 reflector. Communicated by S. Nakano.

A parabolic orbit follows. The comet is probably of short period.

T = 1999 Feb. 20.762 TT  ω = 18.^{o}976
Ω = 92.924 }2000.0
q = 1.58362 AU  i = 29.332
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 18 4^{h}32.^{m}23 −10^{o}24.^{′}2 1.125 2.027 147.^{o}0 15.^{o}4 15.8
23 4 28.24 − 9 36.7 1.075 1.987 148.9 14.9 15.6
28 4 23.69 − 8 33.6 1.031 1.949 150.1 14.6 15.5
Dec. 3 4 18.74 − 7 13.6 0.992 1.912 150.4 14.7 15.3

LEONID METEORS 1998 Numerous further reports received by Central Bureau generally confirm

an extended shower with a preponderance of bright fireballs and a general maximum zenithal hourly rate of perhaps 500. Although impressive, the display was a far cry from the Leonid storms of 1799, 1833 and 1966.

1998 November 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden